Express the negation of the following statements WITHOUT using the negation symbol:
a. ∀x(−2 < x <3)
b. ∀x(0 ≤ x <5)
c. ∃x(−4 ≤ x ≤1)
d. ∃x(−5 < x <−1)
e. ∀x∃y(x2 < y)
1
Expert's answer
2019-09-24T11:18:20-0400


a) "\exists x(x \leq2 \lor x\geq3)"

b) "\exists x(x \lt 0 \lor x\geq 5)"

c) "\forall x(x \lt-4 \lor x\gt1)"

d) "\forall x(x \leq-5 \lor x\geq-1)"

e) "\exists x \forall y (x^2 \geq y)"